Attendees:

Notes:

  • Merlin Fisher-Levine
    • rapid analysis gen3 porting
    • Pair Programming this week on quick-look.
    • Run next week.
    • Web page hasn't found someone to work on.
    • RFCs out for getting spectractor/rapid analysis included.
    • Harvard Ocelot spectrograph 
      • Not running DM ISR so might improve with doing that
      • New student starting soon
      • Resolution around a nm, very roughly
      • Blocking filter to remove blue (610A and above).
    • Getting calibs transferred to the summit.
    • rsync hasn't been installed at the summit - ticket filed.
    • Synchronize repository path names between NCSA and summit.
  • Andrés Alejandro Plazas Malagón
    • Alysha running bias on LATISS, now working for ComCam with less than 10 images.
      • Limitation in queue of 100 total images.
      • Related to DAQ update?
      • Ask Steve to learn if this is resolved.
      • Change in way to take images using callbacks
    • Verification failure due to defects issue
      • Too many cosmic rays.
    • ComCam calibration. Redo DM-28920 there.
      • Transfer to NTS?
    • Bias + Flat script working
    • Friday hardware problems redoing this at the summit
      • Problems this morning as well.
    • K-T and Tim proposed commands for copying commands.
  • Robert Lupton
    • OCPS issues?
    • Imaging survey with LATISS / difference imaging next year.
    • Chilean/commissioning cluster will have a permanent copy of all the data.
      • Software versioning with rebuilding.
  • Meeting this Thursday with Camera Team
    • Bias correction with Camera methods?
      • Not fully converged.
  • Median per row speed-up using random number.